What is the pH of 65 mL of 0.0685 M HNO2?
Answer #1

Consider dissociation of HNO 2 in water as, HNO 2(aq) + H2O (l)phpUIHBnY.png H3O+(aq) +NO 2-(aq)

The expression for dissociation constant is given as, Ka = [H3O+] [NO 2- ]/ [ HNO 2 ] = 7.1 \times 10 -04

To find out equilibrium concentration, let's use ICE table.

Concentration (M) HNO 2 H3O+ NO 2-
Initial 0.0685
Change -X +X +X
Equilibrium 0.0685-X X X

\therefore Ka = (X)(X) / 0.0685 - X = 7.1 \times 10 -04

X 2 / 0.0685 - X = 7.1 \times 10 -04  

X 2 = 7.1 \times 10 -04 (0.0685 - X)

X 2 = 4.86 \times 10 -05 - 7.1 \times 10 -04 X

X 2 + 7.1 \times 10 -04 X - 4.86 \times 10 -05 =0

Compare above equation with a X 2 + b X+ c = 0, we get a = 1, b = 7.1 \times 10 -04 and c= - 4.86 \times 10 -05

Now, solving above equation for X, we get

X = -b +/- \sqrt{}( b 2 - 4 a c ) / 2a

X = -  7.1 \times 10 -04 +/- \sqrt{}(7.1 \times 10 -04) 2 - 4 (1)(- 4.86 \times 10 -05) / 2 (1)

X = -  7.1 \times 10 -04 +/- \sqrt{}5.04 \times 10 -07 + 1.944  \times 10 -04 / 2

X = -  7.1 \times 10 -04 +/- \sqrt{}1.949  \times 10 -04 / 2

X = -  7.1 \times 10 -04 +/- 0.01396 / 2

Therefore, X =-  7.1 \times 10 -04 + 0.01396 / 2   = 6.625 \times 10 -03

Or X = -  7.1 \times 10 -04 - 0.01396 / 2 = - 7.34 \times 10 -03

Acceptable value of X is 6.625 \times 10 -03 .

X = 6.625 \times 10 -03 M = [H3O+] = [NO 2- ]

We have, pH = -log [H3O+] = - log (6.625 \times 10 -03) = 2.18

ANSWER : pH of 0.0685 M HNO 2 solution = 2.18